OpenStudy (anonymous):

The cost of an almirah is Rs 2000.A profit of 10% should be made after a discount of 20% .Find the marked price .(also explain the solution)

OpenStudy (anonymous):

plssssss

OpenStudy (anonymous):

.10Rs=Rs-(.80)(MP)

OpenStudy (anonymous):

solution ??? how u did that ??but r definetely wrong ...the answer is Rs.2750 but I don't know the solution ..

OpenStudy (anonymous):

try this one 1.10Rs=.80MP

OpenStudy (anonymous):

should be \[2750\]

OpenStudy (anonymous):

can u telll how u got the answer @ satelite73

OpenStudy (anonymous):

you want to make at 10% profit, and 10% of 2000 is 200

OpenStudy (anonymous):

you are going to offer a 20% discount, and after the discount you need to get 2200

OpenStudy (anonymous):

so to find the price before the discount take \[2200\div.8=2750\]

OpenStudy (anonymous):

@LILG why an extra 5% profit ??

OpenStudy (lilg132):

to cover for the difference in price because you will be taking out 20% of the final price

OpenStudy (lilg132):

but satellites way is much better then mine

OpenStudy (anonymous):

Let x=retail price and y=marked price x=(1-.20)y-(.10x) 1.10x=.80y y=(1.10/.80)(2000) y=2750

OpenStudy (anonymous):

in other words you need to solve "after a 20% reduction the price is 2200" which is the same as saying "80% of what number is 2200" or in math \[.80\times x =2200\] \[x=2200\div .8\]

OpenStudy (lilg132):

2000 * 0.1 = 200 2000 + 200 = 2200 2200 = 0.8x (To find out how much the price will be to take out 20% discount and still have 10% profit value) 2200 / 0.8 = 0.8x / 0.8 2750 = x

OpenStudy (anonymous):

thnks every1